A 5.5 x10⁴ kg airplane is flying at 85m/s at an altitude of 1.2x10⁴m. Find the plane's



(a) Potential energy ,(b) kinetic energy,and (c) Mechanical energy



Given: m=5.5x10⁴kg



v=85m/s



h=1.2x10⁴m

1
Expert's answer
2022-04-28T16:41:49-0400

(a)

"E_p=-G\frac{MM_e}{(R+h)}"

"E_p=-6.67*10^{-11}\frac{5.5*10^4*5.97*10^{24}}{(6.371*10^6+1.2*10^4)}\\
=-3.4*10^{12}\rm \; J"

(b)

"E_k=\frac{mv^2}{2}=\frac{5.5*10^4*85^2}{2}=2.0*10^8\:\rm J"